# Chain of Strategy Optimization Makes Large Language Models Better Emotional Supporter

### v1

@@ Appendix A LLM Definition and Prompt
We use GPT-4O-mini as the Seeker LLM and employ the prompt shown in fig 5.
Figure 5: Prompt of Seeker LLM
@@
We use GPT-4O-mini as the Supporter LLM and employ the prompt shown in fig 6.
Figure 6: Prompt of Supporter LLM
@@
We use GPT-4O-mini as the Strategy LLM and employ the prompt shown in fig 7.
Figure 7: Prompt of Strategy LLM
@@
We use GPT-4O-mini as the Reward LLM and employ the prompt shown in fig 8. Detailed evaluation criteria are in the appendix B.
Figure 8: Prompt of Reward LLM
@@ C.2 Data Quality Evaluation
We conduct a human evaluation on 100 responses from the ExTES seed dataset, along with 50 preference responses and 50 non-preference responses from ESC-Pro. We consider five metrics (Kang et al., 2024a):
@@
Table 7: Toxicity Analysis Result
@@ C.3 Strategy Analysis
The ExTES dataset originally contains 16 distinct support strategies. To reduce the search space and improve computational efficiency, we merge similar strategies into 8 categories (see Appendix D for details). We analyze the distribution of these strategies across six conversation stages. Given a dialogue with $N$ utterances, each utterance $k$ (where $k=1,…,N$) is assigned to stage $i$ using:
@@ F.1 Dataset Construction Details
When processing the seed data from ExTES, we filtered out entries that did not provide a strategy, contained incorrect strategies, or included the "Others" strategy. The remaining 15 strategies (excluding "Others") used in the ExTES dataset were mapped to the 8 strategies used in our dataset according to a set of predefined rules, which can be found in Table 8.
@@ F.2 Experiments Details
Appendix G Analysis of Different Preference Optimization Algorithms
@@
In addition to preference training based on DPO, we also investigated the performance of several other preference optimization algorithms, including SimPO (Meng et al., 2024), IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024), and ORPO (Hong et al., 2024). We tested these algorithms in both LoRA and full fine-tuning settings, maintaining the training parameters consistent with DPO, with adjustments made only to algorithm-specific parameters. For instance, the hyperparameter $\beta$ was adjusted to suit each algorithm, and algorithm-specific parameters, such as GEMMA for SimPO, chosen weight and rejected weight for KTO, were tuned accordingly. A detailed list of the hyperparameters used for each algorithm can be found in the Table 13 and Table 14.
For all algorithms except KTO, we used the ESC-Pro dataset as the training set, which contains preference pairs directly derived from the original ESC-Pro data. In the case of KTO, the training set was derived by splitting each preference pair in the ESC-Pro dataset and removing duplicates. The experimental results are shown in Table 9, Table 10 and Table 11. From the results, it is evident that all tested preference optimization algorithms performed effectively after training on the ESC-Pro dataset, with some methods achieving higher performance than DPO. These results further validate the efficacy and versatility of the ESC-Pro dataset for optimizing preference-based dialogue strategies.
@@ Appendix H Case Study
Appendix I Manual Evaluation Implementation Details
We conducted manual evaluations and outsourced the evaluation tasks to anonymous crowd workers. Detailed guidelines and instructions were provided for each rating criterion to ensure the accuracy and consistency of the evaluations, as outlined in appendix C.2. For each evaluation task, we required three evaluators to assess 100 samples based on the corresponding metrics.

### v3

@@ Appendix A LLM Definition and Prompt
We use GPT-4o-mini as the Seeker LLM and employ the prompt shown in Figure 5.
Figure 5: Prompt of seeker LLM.
@@
We use GPT-4o-mini as the Supporter LLM and employ the prompt shown in Figure 6.
Figure 6: Prompt of supporter LLM.
@@
We use GPT-4o-mini as the Strategy LLM and employ the prompt shown in Figure 7.
Figure 7: Prompt of strategy LLM.
@@
We use GPT-4o-mini as the Reward LLM and employ the prompt shown in Figure 8. Detailed evaluation criteria are in the Appendix B.
Figure 8: Prompt of reward LLM.
@@ C.2 Data Quality Evaluation
We conduct a human evaluation on 100 responses from the ExTES seed dataset, along with 50 preference responses and 50 non-preference responses from ESC-Pro. We consider five metrics (Kang et al., 2024):
@@
Table 7: Toxicity analysis of different results.
@@ C.3 Strategy Analysis
The ExTES dataset originally contains 16 distinct support strategies. To reduce the search space and improve computational efficiency, we merge similar strategies into 8 categories (see Appendix D for details). We analyze the distribution of these strategies across six conversation stages. Given a dialogue with $N$ utterances, each utterance $k$ (where k=1,,Nk=1,…,N) is assigned to stage $i$ using:
@@ Appendix D Definitions of Strategies
The reduction from 16 to 8 strategies is a deliberate design decision based on both empirical observations and theoretical coherence. Many of the original strategies in Zheng et al. (2024) have significant semantic overlap and are frequently confused by annotators in preliminary labeling exercises. Detailed categories are shown in Table 8.
For example: Reflective Statements, Clarification, Normalize Experiences, and Emotional Validation all aim to validate or mirror the user’s feelings. These are consistently grouped together under the broader category of Emotional Validation by human annotators. Similarly, Suggest Options, Collaborative Planning, Stress Management, and Promote Self-Care Practices are all oriented toward joint problem-solving, and are thus unified under Collaborative Planning.
We make these groupings with the dual goal of: (1) Improving label consistency and inter-annotator agreement (IAA), and (2) Focusing the learning signal on distinct and meaningful strategic differences rather than fine-grained variations that models (and even humans) struggle to reliably differentiate.
@@ F.1 Dataset Construction Details
When processing the seed data from ExTES, we filtered out entries that did not provide a strategy, contained incorrect strategies, or included the “Others” strategy. The remaining 15 strategies (excluding “Others”) used in the ExTES dataset were mapped to the 8 strategies used in our dataset according to a set of predefined rules, which can be found in Table 8.
@@ F.2 Experimental Details
Appendix G Additional Experimental Results
G.1 Analysis of Different Preference Optimization Algorithms
@@
In addition to preference training based on DPO, we also investigated the performance of several other preference optimization algorithms, including SimPO (Meng et al., 2024), IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024), and ORPO (Hong et al., 2024). We tested these algorithms in both LoRA and full fine-tuning settings, maintaining the training parameters consistent with DPO, with adjustments made only to algorithm-specific parameters. For instance, the hyperparameter $\beta$ was adjusted to suit each algorithm, and algorithm-specific parameters, such as Gemma-2 for SimPO, chosen weight and rejected weight for KTO, were tuned accordingly. A detailed list of the hyperparameters used for each algorithm can be found in the Table 14 and Table 15.
For all algorithms except KTO, we used the ESC-Pro dataset as the training set, which contains preference pairs directly derived from the original ESC-Pro data. In the case of KTO, the training set was derived by splitting each preference pair in the ESC-Pro dataset and removing duplicates. The experimental results are shown in Table 9, Table 10 and Table 11. From the results, it is evident that all tested preference optimization algorithms performed effectively after training on the ESC-Pro dataset, with some methods achieving higher performance than DPO. These results validate the efficacy and versatility of the ESC-Pro dataset for optimizing preference-based dialogue strategies.
[…]
Qwen2.5-32B-Instruct 37.37 1.40 41.97 24.37
Direct-Refine 34.87 1.68 40.69 23.14
Self-Refine 14.63 0.98 26.97 21.74
w/ Example 20.66 2.65 24.40 21.93
SFT-LoRA 37.69 1.60 42.09 24.40
CSO-LoRA 38.53 1.29 43.95 24.65
Table 12: Performance comparison on Qwen2.5-32B-Instruct using LoRA-based preference optimization.
G.2 Results of Large-Scale Backbone
We conduct additional experiments on Qwen2.5-32B-Instruct using LoRA-based preference optimization. As shown in Table 12, CSO-LoRA consistently outperforms all baselines across all four metrics, and long-term user satisfaction (R-L). Compared to standard supervised fine-tuning (SFT-LoRA), CSO-LoRA improves strategy quality by +0.84 and reduces strategy bias while further enhancing user-centered metrics. These results confirm that CSO remains effective even at larger scales, demonstrating strong scalability and robustness in enhancing both adaptability and emotional intelligence in LLM-based emotional support.
@@ Appendix H Case Study
In Table 17, we provide a case study to demonstrate that CSO can implicitly preserve emotional continuity over multiple turns by selecting context-appropriate strategies. This case shows that CSO can implicitly maintain emotional consistency across turns by adapting strategies to the seeker’s changing emotions. The supporter begins with validation and empathy, then gradually introduces more proactive strategies like planning and reframing as the seeker becomes more hopeful. This smooth progression show that CSO responds not just to the current emotion, but also supports the emotional flow of the conversation as a whole.
@@
Appendix I Human Evaluation Details
We conducted human evaluations and outsourced the evaluation tasks to anonymous crowd workers. We employed three evaluators for the human evaluation process, all of whom had prior annotation experience to ensure reliability. Before conducting evaluations, they received targeted training to thoroughly understand the criteria for each metric, promoting consistency and accuracy in their judgments. Evaluators were compensated based on the number of annotated samples they completed. To assess annotation consistency, we report the inter-annotator agreement using Cohen’s Kappa scores, which are provided in the final column of Tables 1 and 3. Detailed guidelines and instructions were provided for each rating criterion to ensure the accuracy and consistency of the evaluations, as outlined in appendix C.2. For each evaluation task, we required three evaluators to assess 100 samples based on the corresponding metrics.
Conversation Background
Scene Depression and Low Mood
Description I’ve been feeling emotionally exhausted and overwhelmed by everything happening in my life. It’s hard to stay positive when it feels like things keep piling up and nothing changes.
Role Emotion / Strategy Utterance
Seeker Sadness I’ve been feeling really overwhelmed lately. I don’t know how to handle everything that’s going on.
Supporter Emotional Validation It sounds like you’re carrying a lot right now, and it’s understandable that you’re feeling overwhelmed. It’s okay to feel that way.
Seeker Frustration Yeah, it’s just so much, and I feel like I can’t get a grip on it.
Supporter Empathetic Statements It’s tough when everything seems like it’s piling up. I can hear how much you’re struggling with it.
Seeker Hopelessness I’m trying to stay positive, but it’s hard.
Supporter Affirmation It’s really impressive that you’re trying to stay positive despite everything. That shows real strength.
Seeker Overwhelm I just don’t know where to start.
Supporter Collaborative Planning Let’s take it step by step. What’s one small thing you could focus on right now?
Seeker Hope Maybe I can tackle some smaller tasks to get the ball rolling.
Supporter Provide Different Perspectives That’s a great perspective. Starting small can help you feel more in control and build momentum.
Seeker Defeat I feel like I’ve been stuck in this cycle, and nothing changes.
Supporter Reframe Negative Thoughts It’s easy to feel stuck when it seems like things aren’t improving. But every small step forward counts, even if it doesn’t feel like it.
Seeker Acceptance You’re right. Maybe I need to be kinder to myself through this.
Supporter Avoid Judgment and Criticism Absolutely. Being kind to yourself is key. You’re doing your best, and that’s enough.
Seeker Hopeful I think I can manage it now, one step at a time.
Supporter Emotional Validation That’s a wonderful mindset. Taking it one step at a time will help you regain your balance.
Table 17: Dialogue turns with emotional Context and support strategies